Container A holds 782 mL of ideal gas at 2.80 atm. Container B holds 174 mL of ideal gas at 4.40 atm. If the gases are allowed to mix together, what is the resulting pressure?
1
Expert's answer
2018-12-07T02:14:37-0500

Answer on Question #83513

Container A holds 782 mL of ideal gas at 2.80 atm. Container B holds 174 mL of ideal gas at 4.40 atm. If the gases are allowed to mix together, what is the resulting pressure?

Solution

We should use Dalton's Law of Partial Pressures to calculate the resulting pressure:

Ptotal = PA + PB

Find pressure (PA) of ideal gas that was in the container A when two gases were mixed together. We should use Boyle's Law:

P1V1 = P2V2

where P1 = 2.80 atm, V1 = 782 mL, P2 = PA , V2 = 782 mL + 174 mL = 956 mL

2.80*782 = PA*956

PA = 2.29 atm

Find pressure (PB) of ideal gas that was in the container B when two gases were mixed together. We should use Boyle's Law:

P1V1 = P2V2

where P1 = 4.40 atm, V1 = 174 mL, P2 = PB , V2 = 782 mL + 174 mL = 956 mL

4.40*174 = PB*956

PB = 0.8 atm

Ptotal = PA + PB= 2.29 atm + 0.8 atm = 3.09 atm

Answer: 3.09 atm